Hi there,
I'm not exactly sure how to set up this problem. Any help is very much appreciated!
Let f(x) = 3x^2 + 7x – 5 and g(x) = 2 – x.
(a) Evaluate the function g – f for x = –1. That is, find (g – f)(–1).

(b) Evaluate the function fg for x = –1. That is, find (f g)( –1).


(c) Find the difference function (f – g)(x) and simplify the results.

Answer by ikleyn(53751)

.
Hi there,
I'm not exactly sure how to set up this problem. Any help is very much appreciated!
Let f(x) = 3x^2 + 7x – 5 and g(x) = 2 – x.

(a) Evaluate the function g – f  for x = –1. That is, find (g – f)(–1).    <<<--- Calculate f(-1) and g(-1), then subtract.


(b) Evaluate the function fg for x = –1. That is, find (f g)( –1).         <<<--- Calculate f(-1) and g(-1), then multiply.



(c) Find the difference function (f  – g)(x) and simplify the results.     <<<--- Calculate the difference of the polynomials.
                                                                                  It is  (3x^2 + 7x – 5) - (2 – x) = 3x^2 + 8x -7.